52% of the employees in a business are women. 452 employees are men. How many employees are women?
Found 2 solutions by rfer, dhritiman:
Answer by rfer(16322)

452/0.48=942 total
942*0.52=490 women

Answer by dhritiman(2)

52% women so 100-52=48% men
hence 48x/100=452 [let x be total ]
and from here we can have x value
then to find the women members
452/48*52=1469/3=489.66 equivalent to 490.